# Fractional phase jumps in stochastic systems with tilted periodic double-well potentials

### v1

@@ Abstract
We present a theoretical investigation of the stochastic dynamics of a damped particle in a tilted periodic potential with a double well per period. By applying the matrix continued fraction technique to the Fokker-Planck equation in conjunction with the full counting statistics and master equation approaches, we determine the rates of specific processes contributing to the system’s overall dynamics. At low temperatures, the system can exhibit one running state and two distinct locked metastable states. We focus primarily on two aspects: the dynamics of phase jumps, which are rare thermally induced particle jumps over potential maxima, and their impact on the overall velocity noise; and the retrapping process, involving the transition from the running to the locked metastable states. We demonstrate the existence of fractional (in units of $2\pi$) phase slips that differ qualitatively from conventional $2\pi$ jumps observed in single-well systems. Fractional phase slips significantly influence the system dynamics even in regimes dominated by dichotomous-like switching between running and locked states. Furthermore, we introduce a simple master equation approach that proves effective in analyzing various stages of the retrapping process. Interestingly, our analysis shows that even for a system featuring a well-developed double-well periodic potential, there exists a broad parameter range where the stochastic dynamics can be accurately described by an effective single-well periodic model. The techniques introduced here allow for valuable insights into the complex behavior of the system, offering avenues for understanding and controlling its steady-state and transient dynamics, which go beyond or can be complementary to direct stochastic simulations.
@@ I Introduction
The stochastic dynamic motion of a particle in tilted periodic potentials plays an important role in the studies of various physical phenomena including Josephson junctions (JJs) [1, 2, 3, 4, 5], microparticles confined in shaped laser beams [6, 7, 8, 9, 10, 11, 12, 13], dynamics of charge density waves [14], crystal surface melting [15, 16], ratchet and molecular motors [17, 18, 19], cold atoms in optical lattices [20], different bio-physical processes [21] as well as in investigations of phenomena such as anomalous diffusion and memory effects [22, 23, 24, 25, 26]. A well-paid effort has already been invested in the theoretical analysis of such systems [27, 28, 29, 30], yet there are still plenty of open questions often inspired by the recent experimental realizations in different systems, e.g., Refs. [4, 13, 31, 32]. Moreover, experimental progress in different fields called for readdressing some old problems such as escape and retrapping of the Brownian particle from or to potential minima [33, 34, 35, 36, 37, 24], the statistics of thermally activated jumps of the particle by integer multiples of $2\pi$ (i.e., phase jumps) [38, 39, 40, 35] and multistability [24, 41, 42].
@@
where $i_{b}$ is a static external bias force, i.e., a potential tilt, $U_{0}(\varphi)$ is the untilted periodic double-well potential, where the coefficient $\alpha$ tunes the ratio of the first to the second harmonic contribution and $\varphi=2\pi x/\lambda$, where $x$ is the coordinate (position) and $\lambda$ the period. Consistent with papers focusing, for example, on the Josephson junctions, we call the variable $\varphi$ phase. The potential (1) plays an important role in the theoretical description of numerous physical systems, including JJs with substantial second harmonic in the current phase relation (CPR) [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 4, 55, 56] in particular in Josephson diodes [57, 58, 31] or various ratchet systems [59, 60, 61] and molecular motors [62, 63].
@@
A careful analysis of the motion of the particle in the tilted double-well potential has already led to some important results. For example, it explained the existence of two critical escape currents from the superconducting to the resistive state observed experimentally for the JJs with doubly degenerate ground state (i.e., $\varphi$-Junctions) [48], pointed to rather nontrivial retrapping dependencies of the phase which can lead to a butterfly effect [50] and showed the existence of chaotic phase trajectories in various generalizations of the famous RCSJ model [64, 65]. The case $\alpha=0$, where only the second harmonic is present in the potential, plays an important role in the studies of unconventional junctions which can undergo the so-called 00-$\pi$ transition [66, 54]. Moreover, a Brownian particle moving in the potential (1) with additional harmonic terms, characterized by asymmetric mobility considering bias force, is an archetypal model of the ratchet and diode systems [59, 61, 57, 58].
@@
A crucial component in all of the above phenomena is noise. In combination with different dampings, a system with potential (1) can show a wide variety of regimes, each important for different physical realizations. Here, we provide a systematic analysis of a general case. We start with the simple strong damping parameter regime and proceed to the more complicated intermediate and weak damping regimes. We use velocity-noise to identify three main dynamic regimes [35]: the thermal noise regime; the phase-jumps (PJs) regime, and the switching regime. In each, we focus on the analysis of the dominant dynamics. In particular, we investigate the statistics of the phase jumps and their contribution to the overall velocity noise in the PJs regime and the escape and retrapping processes in the switching one. For this purpose, we combine the matrix continued-fraction (MCF) method with other techniques. In particular, in Sec. II we introduce a combination of the matrix continued-fraction (MCF) technique [27] with full counting statistics (FCS) [67, 68, 69] for the analysis of the multiple PJs. Its biggest advantage over the stochastic simulations is that it is straightforward to access the steady state. As such, it is suitable for the calculation of rates for rare events. Therefore, in the relevant regime, this method allows decomposition of the phase dynamics into independent elementary processes [70, 71, 72], constituted by single or multiple phase jumps. Using this method, we demonstrate the existence of the so-called fractional PJ in Sec. III.2.1. In Sec. III.2.2 we investigate retrapping processes in the switching regime. Here we combine the MCF with an effective master equation approach describing the transition of the system between its three metastable regimes. On top of the steady-state studies, we also discuss in Sec. III.2.3 a dynamical retrapping scenario.
@@ II Model and Methods
Figure 1: (a) Examples of $v-i_{b}$ characteristics representing the strong damping ($\gamma=5$ plotted with red lines) and the weak damping regime ($\gamma=0.2$ plotted with black lines). The thin dashed lines show the deterministic (noise-free) solutions. The solid thick lines represent stationary solutions for temperature $\Theta=0.1$. (b) -(c) Illustrations of the potential Eq. (1) with $\Gamma^{A}_{n}$ and $\Gamma^{B}_{n}$ being the rates of the forward / backward phase jumps over $n$ local maxima. In general the distance $d$ is not equal to $\pi$.
@@ II.1 Model and dynamical regimes
We consider a stochastic motion of a particle in a potential $U(\varphi)$ described by the dimensionless Langevine equations
@@
where $v$ is the velocity of the particle, $\gamma$ is the friction coefficient and $\zeta$ represents a Gaussian white noise with the zero mean $\langle\zeta(\tau)\rangle=0$ and correlation function $\langle\zeta(\tau_{1})\zeta(\tau_{2})\rangle=2\gamma\Theta\delta(\tau_{1}-\tau_{2})$ where $\Theta$ is the dimensionless temperature. The associated Fokker-Planck equation [27] of Eqs. (2) for the probability distribution function $W(\varphi,v,\tau)$ in the case of potential (1) reads
@@
follows closely the standard MCF method [27, Sec. 11.5] and the derivation of (zero-frequency) velocity noise
@@
Figs. 1(b)-(c) illustrate the system for zero and finite bias force. The potential (1) has for $\alpha<2$ and zero bias two types of minima marked as $A$ and $B$ and two types of maxima ($a$ and $b$). Consequently, there are also two critical bias forces: $i_{cA}$ where the minimum $A$ and maximum $a$ disappear and $i_{cB}$ where the minimum $B$ and maximum $b$ disappear. The system can be in three distinct stationary regimes. Namely, the particle can be running, i.e., at high bias force, or it can be locked in one of the potential minima types. Consequently, for the noiseless case, the initial position of the particle can, depending on the damping, play an important role even for the steady state, as illustrated in Fig. 1(a).
@@
In contrast, there are two possible escape tilts for weak damping ($\gamma=0.2$). If the particle initially is trapped at the minimum $A$, it will obtain enough inertia to overcome the still existing maximum $b$ already at the first critical force $i_{cA}$. However, if it is initially trapped at the minimum $B$ it will stay there until the second critical force $i_{cB}$ is reached. The retrapping is also more complicated than in the strong-damping case. If the particle is already running, then it has enough inertia to overcome the local maxima existing below the critical force $i_{cB}$ or even $i_{cA}$ up to the actual retrapping force $i_{r}$ [73], black dashed lines in Fig. 1(a). The retrapping force is determined by the energy balance between the energy supply of the bias force and the dissipation [27, 50]. This means that there is a region of coexistence of the running and the locked state solutions. The question in which minimum will the particle be trapped requires careful analysis because it is a parameter-sensitive process which can lead to a deterministic butterfly effect [50, 74].
@@
There are three main components that contribute to the overall velocity noise of this model [33, 35]. The thermal noise component dominates close to the equilibrium $i_{b}=0$. The second component is the switching noise coming from the switching between running and locked states, which is, for low enough temperatures, an effective dichotomous-like process. The third component is the shot noise (PJs regime) related to rare jumps of the particle over single or multiple maxima. The regimes where particular components prevail can be identified from the Fano factor $F\equiv S/(2\pi\left\langle v\right\rangle)$ [35], which is a normalized noise-to-signal ratio.
II.2 Full-counting statistics for the phase jump dynamics
The phase jumps are, for low enough temperature and weak bias force, well defined distinct (rare) events that significantly influence the overall dynamics. To calculate the rates of these events, we have adapted the full-counting statistics technique previously used to study jump probabilities in single-harmonic systems [67, 68, 69, 35]. The double-well character of the potential (1) requires some generalizations of this method which we present here.
In the first step we have approximated the solution of the Fokker-Planck equation (II.1) for sufficiently low biases and temperatures by a weighted sum of quasi-equilibrated sharp ($\Theta\ll 1$) Gaussian distributions [69] around the two types od local minima
@@ II.2 Full-counting statistics for the phase jump dynamics
Here $\varphi^{A}_{m}$ and $\varphi^{B}_{m}$ are the positions of the $m$-th potential minima and $P^{A}_{m}(\tau)$ and $P^{B}_{m}(\tau)$ are the corresponding time-dependent weights. These are assumed to satisfy the (Markovian) master equations (ME)
$\frac{dP^{A}_{m}}{d\tau}$ $=$ $\sum_{n}\left(\Gamma^{A}_{2n}P^{A}_{m-n}-\Gamma^{A}_{n}P^{A}_{m}+\Gamma^{B}_{2n-1}P^{B}_{m-n}\right),$ (7)
$\frac{dP^{B}_{m}}{d\tau}$ $=$ $\sum_{n}\left(\Gamma^{B}_{2n}P^{B}_{m-n}-\Gamma^{B}_{n}P^{B}_{m}+\Gamma^{A}_{2n+1}P^{A}_{m-n}\right),$ (8)
where $\Gamma^{A}_{n}$ is the rate of a phase jump from the potential well $A$ and $\Gamma^{B}_{n}$ from the well $B$ over $n$ potential local maxima to another local minimum. Rates with even $n$ belong to phase jumps between minima of the same kind ($A\rightarrow A$, $B\rightarrow B$) and odd ones to phase jumps between minima of different kinds ($A\rightarrow B$, $B\rightarrow A$). The negative $n$’s correspond to the jump rates in the direction opposite the slope of the bias (up the hill). Following the standard FCS methodology [69], we can evaluate the $k$-dependent cumulant generating function (CGF) for long times from the ME and equate it with the CGF of the full model.
@@
$\mathbf{H}^{\textrm{PJ}}(k)=\left(\begin{array}[]{cc}-\sum\limits_{n}\Gamma^{A}_{n}+\Gamma^{A}_{2n}e^{ik2\pi n}&\sum\limits_{n}\Gamma^{B}_{2n-1}e^{ik(2\pi n-d)}\\ \sum\limits_{n}\Gamma^{A}_{2n+1}e^{ik(2\pi n+d)}&-\sum\limits_{n}\Gamma^{B}_{n}+\Gamma^{B}_{2n}e^{ik2\pi n}\end{array}\right).$ (20)
The $d$ factor in the exponents of above equations is the distance between the neighboring minima $d=\varphi^{B}_{m}-\varphi^{A}_{m}$ [see Fig. 1(c)] which depends on the potential parameters $\alpha$ and $i_{b}$. In contrast to even phase jumps, where the distance traveled is always an integer multiple of $2\pi$, odd jumps overcome a distance of $2\pi n+d$ where in general $d\neq\pi$, therefore, we call these jumps fractional.
Analogously to Appendix in Ref. [35] we use MCF to calculate the two eigenvalues $\lambda_{0}(k),\>\lambda_{1}(k)$ with the largest real parts and related eigenvectors $u_{0}(k,\varphi,v)$, $u_{1}(k,\varphi,v)$  [27, Sec. 9.3]. The above two eigenvalues are in the relevant regime well separated from all the others.
@@
and $\mathbf{U}(k)$ is the square matrix of vectors $U_{0}(k)$ and $U_{1}(k)$. Note that due to the second harmonics in the potential and similarly to other non-Hermitian Hamiltonian systems [75, 76], the eigenvalues (22) can have complicated topological properties, as shown in Fig. 2. The two eigenvalues can, for small enough $\alpha$, connect at $k_{c}=1/2+m$, where $m$ is an integer, and smoothly continue each other on the Riemann surface; see Fig. 2(a). Therefore, the real parts of these two eigenvalues, plotted as functions of $k$, touch on $k_{c}$ [Fig. 2(b)], and there is a discontinuity in the imaginary part of the eigenvalues at the same points [Fig. 2(c)]. These discontinuities must be treated with care in the numerical evaluation of the eigenvalues and eigenvectors.
Figure 2: The two eigenvalues with the biggest real part plotted at the Riemann surface (a) and their projections to the real (b) and the imaginary (c) plane.
Having the reconstructed matrix $\mathbf{H}^{\textrm{PJ}}$ one can evaluate the rates of the even phase jumps ($2\pi n$ PJ between the same kind of minima) using the transformations
@@
and the odd ones between the different minima types as
@@
$\langle v\rangle$ $=\lim_{\tau\rightarrow\infty}-\frac{i}{\tau}\left.\frac{\partial\mathcal{F}^{\textrm{PJ}}(k,\tau)}{\partial k}\right|_{k=0},$ (25)

### v2

# Fractional positional jumps in stochastic systems with tilted periodic double-well potentials

@@ Abstract
We present a theoretical investigation of the stochastic dynamics of a damped particle in a tilted periodic potential with a double well per period. By applying the matrix continued fraction technique to the Fokker-Planck equation in conjunction with the full counting statistics and master equation approaches, we determine the rates of specific processes contributing to the system’s overall dynamics. At low temperatures, the system can exhibit one running state and two distinct locked metastable states. We focus primarily on two aspects: the dynamics of positional jumps, which are rare thermally induced particle jumps over potential maxima, and their impact on the overall velocity noise; and the retrapping process, involving the transition from the running to the locked metastable states. We demonstrate the existence of fractional (in units of $2\pi$) positional slips that differ qualitatively from conventional $2\pi$ jumps observed in single-well systems. Fractional positional slips significantly influence the system dynamics even in regimes dominated by dichotomous-like switching between running and locked states. Furthermore, we introduce a simple master equation approach that proves effective in analyzing various stages of the retrapping process. Interestingly, our analysis shows that even for a system featuring a well-developed double-well periodic potential, there exists a broad parameter range where the stochastic dynamics can be accurately described by an effective single-well periodic model. The techniques introduced here allow for valuable insights into the complex behavior of the system, offering avenues for understanding and controlling its steady-state and transient dynamics, which go beyond or can be complementary to direct stochastic simulations.
@@ I Introduction
The stochastic dynamic motion of a particle in tilted periodic potentials plays an important role in the studies of various physical phenomena including Josephson junctions (JJs) [1, 2, 3, 4, 5], microparticles confined in shaped laser beams [6, 7, 8, 9, 10, 11, 12, 13], dynamics of charge density waves [14], crystal surface melting [15, 16], ratchet and molecular motors [17, 18, 19], cold atoms in optical lattices [20], and different biophysical processes [21] as well as in investigations of phenomena such as anomalous diffusion and memory effects [22, 23, 24, 25, 26]. A well-paid effort has already been invested in the theoretical analysis of such systems [27, 28, 29, 30], yet there are still plenty of open questions often inspired by the recent experimental realizations in different systems, e.g., Refs. [4, 13, 31, 32]. Moreover, experimental progress in different fields called for readdressing some old problems such as escape and retrapping of the Brownian particle from or to potential minima [33, 34, 35, 36, 37, 24], the statistics of thermally activated jumps of the particle by integer multiples of $2\pi$ [38, 39, 40, 35] and multistability [24, 41, 42].
@@
where $i_{b}$ is a static external bias force, i.e., a potential tilt, $U_{0}(\varphi)$ is the untilted periodic double-well potential, where the coefficient $\alpha$ tunes the ratio of the first to the second harmonic contribution and $\varphi$ is the dimensionless position. The potential (1) plays an important role in the theoretical description of numerous physical systems, including JJs with substantial second harmonic in the current phase relation (CPR) [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 4, 55, 56] in particular in Josephson diodes [57, 58, 31, 59, 60] (where the variable $\varphi$ is the superconducting phase difference, i.e., the Josephson phase), or various ratchet systems [61, 62, 63], molecular motors [64, 65] and multiband superconductors [66].
@@
A careful analysis of the motion of the particle in the tilted double-well potential has already led to some important results. For example, it explained the existence of two critical escape currents from the superconducting to the resistive state observed experimentally for the JJs with doubly degenerate ground state (i.e., $\varphi$ junctions) [48], pointed to rather nontrivial retrapping dependencies of the Josephson phase which can lead to a butterfly effect [50] and showed the existence of chaotic Josephson phase trajectories in various generalizations of the famous RCSJ model [67, 68]. The case $\alpha=0$, where only the second harmonic is present in the potential, plays an important role in the studies of unconventional junctions which can undergo the so-called 00-$\pi$ transition [69, 54]. Moreover, a Brownian particle moving in the potential (1) with additional harmonic terms, characterized by asymmetric mobility considering the bias force, is an archetypal model of the ratchet and diode systems [61, 63, 57, 58].
@@
A crucial component in all of the above phenomena is noise. In combination with different dampings, a system with potential (1) can show a wide variety of regimes, each important for different physical realizations. Here, we provide a systematic analysis of a general case. We start with the simple strong-damping parameter regime and proceed to the more complicated intermediate- and weak-damping regimes. We use velocity noise to identify three main dynamic regimes [35]: the thermal-noise regime; the positional-jump (PJ) regime, and the switching regime. In each, we focus on the analysis of the dominant dynamics. In particular, we investigate the statistics of the positional jumps and their contribution to the overall velocity noise in the PJ regime and the escape and retrapping processes in the switching one. For this purpose, we combine the matrix continued-fraction (MCF) method with other techniques. In particular, in Sec. II we introduce a combination of the MCF technique [27] with full counting statistics (FCS) [70, 71, 72] for the analysis of the multiple PJs. Its biggest advantage over stochastic simulations is the straightforward access to the steady state. As such, it is suitable for the calculation of rates for rare events. Therefore, in the relevant regime, this method allows decomposition of the particle dynamics into independent elementary processes [73, 74, 75], constituted by single or multiple positional jumps. Using this method, we demonstrate the existence of the so-called fractional PJs in Sec. III.2.1. In Sec. III.2.2 we investigate retrapping processes in the switching regime. Here we combine the MCF with an effective master equation approach describing the transition of the system between its three metastable regimes. On top of the steady-state studies, we also discuss in Sec. III.2.3 a dynamical retrapping scenario.
@@ II Model and Methods
Figure 1: (a) Examples of $v-i_{b}$ characteristics representing the strong damping ($\gamma=5$) and the weak damping ($\gamma=0.2$) regimes. The solid thick lines show stationary solutions for temperature $\Theta=0.1$. The thin dashed lines show the deterministic (noise-free) solutions. The weak damping case is sensitive to the initial conditions. The black dashed line shows a scenario in which we start with a large force $i_{b}$, and therefore finite velocity $v$, and then lower $i_{b}$ slowly enough to reach steady state at each point. The opposite scenario, where we start with $i_{b}=0$ is represented by a blue dashed line for $\varphi_{0}$ initially locked in the minimum $A$ and red dashed line for the minimum $B$. (b) -(c) Illustrations of the potential Eq. (1) with $\Gamma^{A}_{n}$ and $\Gamma^{B}_{n}$ being the rates of the forward / backward positional jumps over $n$ local maxima. (d1)-(d6) Potential profiles for values of $\alpha$ used in the paper and four bias forces $i_{b}=0$, $0.3$, $0.6$ and $0.9$. The vertical gray lines at $\varphi=\pm\pi/2$ are guides for the eyes that highlight that the minima do not sit at these values and the minima distance $d$ is not necessarily equal to $\pi$.
@@ II.1 Model and dynamical regimes
We consider a stochastic motion of a particle in a potential $U(\varphi)$ with $\alpha\geq 0$ described by the dimensionless Langevin equations
@@
where $v$ is the dimensionless velocity of the particle, $\gamma$ is the friction coefficient and $\zeta$ represents a Gaussian white noise with the zero mean $\langle\zeta(\tau)\rangle=0$ and correlation function $\langle\zeta(\tau_{1})\zeta(\tau_{2})\rangle=2\gamma\Theta\delta(\tau_{1}-\tau_{2})$ where $\Theta$ is the dimensionless temperature. In what follows, we assume that the system is in the classical regime. This means that $\Theta$ is much higher than some critical $\Theta_{c}$ of the quantum-classical transition. This critical temperature depends on details of the physical realization of the system and is not part of the effective model in Eq. (2) which already assumes $\Theta\gg\Theta_{c}$. This assumption must be verified for each particular realization of the problem because quantum effects, such as macroscopic quantum tunneling known from Josephson junctions [76, 77, 78, 4], can significantly alter the particle dynamics if this condition is not fulfilled.
The associated Fokker-Planck equation [27] of Eqs. (2) for the probability distribution function $W(\varphi,v,\tau)$ in the case of potential (1) reads
@@
follows closely the standard MCF method (see Ref. [27, Sec. 9] for a general introduction to MCF and Ref. [27, Sec. 11.5] for the variant here used) and the derivation of (zero-frequency) velocity noise
@@
Figs. 1(b)-(c) illustrate the system for zero and finite bias force, and additional examples relevant for our study are plotted in panels (d1)-(d6). The potential (1) has for $\alpha<2$ and zero bias two types of minima marked as $A$ and $B$ and two types of maxima marked as $a$ and $b$. Consequently, there are also two critical bias forces: $i_{cA}$ where the minimum $A$ and maximum $a$ disappear and $i_{cB}$ where the minimum $B$ and maximum $b$ disappear. The system can be in three distinct stationary regimes. Namely, the particle can be running, i.e., at high bias force, or it can be locked in one of the potential minima types. Consequently, for the noiseless case, the initial position of the particle can, depending on the damping, play an important role even for the steady state, as illustrated in Fig. 1(a).
@@
In contrast, there are two possible escape tilts for weak damping ($\gamma=0.2$). If the particle initially is trapped at the minimum $A$, it will obtain enough inertia to overcome the still existing maximum $b$ already at the first critical force $i_{cA}$. However, if it is initially trapped at the minimum $B$ it will stay there until the second critical force $i_{cB}$ is reached. The retrapping is also more complicated than in the strong-damping case. If the particle is already running, then it has enough inertia to overcome the local maxima existing below the critical force $i_{cB}$ or even $i_{cA}$ up to the actual retrapping force $i_{r}$ [79], black dashed lines in Fig. 1(a). The retrapping force is determined by the energy balance between the energy supply of the bias force and the dissipation [27, 50]. This means that there is a region of coexistence of the running and the locked state solutions.
The question of in which minimum the particle will be trapped requires careful analysis because it is a parameter-sensitive process. We illustrate this in Fig. 2 using three sets of parameters for which all three solutions can be realized depending on the initial conditions, with the force $i_{b}$ being close to the retrapping one. The top row shows examples of separatrix curves and the respective regions of initial conditions that lead to running (white), locked $A$ (blue), and locked $B$ (red) solutions. Separatrices are $2\pi$ periodic and become more complex with decreasing $\gamma$ and $i_{b}$. As a consequence, the retrapping trajectory of a particle can be very complicated. This is illustrated in the bottom row of Fig. 2. The initial conditions are marked by empty symbols (circles or triangles), and the final state is marked by filled ones. The right column illustrates a setting in which even a tiny monotonic change in the initial position of the particle can lead to switching between the two locked final states, that is, a deterministic butterfly effect discussed in more detail, e.g., in Refs. [50, 80].
Figure 2: Top row: Separatrices (black lines) shown for different dampings and bias forces that illustrate the complicated retrapping dynamics of the noiseless case. Any particle within the white region will keep running forever, particles within the blue and red regions will get trapped following paths bounded by the regions of the same color as their initial state (they cannot cross a separatrix). For example, a motionless particle placed initially close to the top of the higher potential maximum, but just right of the running-locked separatrix as marked by the green empty circles ($v=0,$ $\varphi=\varphi^{b}+\delta$), will end in a minimum marked by the green filled circle. Bottom row: Examples of the retrapping trajectories for initial conditions marked by the empty circles and triangles. The blue lines mark the trajectories bounded to the locked solution $A$ and the red lines to the locked solution $B$.
@@
There are three main components that contribute to the overall velocity noise of this model [33, 35]. The thermal noise component dominates close to the equilibrium $i_{b}=0$. The second component is the switching noise coming from the switching between running and locked states, which is, for low enough temperatures, an effective dichotomous-like process. The third component is the shot noise (PJ regime) related to rare jumps of the particle over single or multiple maxima. The regimes where particular components prevail can be identified from the Fano factor $F\equiv S/(2\pi\left\langle v\right\rangle)$ [35], which is a normalized noise-to-signal ratio.
II.2 Full-counting statistics for the positional jump dynamics
The positional jumps are, for low enough temperature and weak bias force, well defined distinct (rare) events that significantly influence the overall dynamics. To calculate the rates of these events, we have adapted the full-counting statistics technique previously used to study jump probabilities in single-harmonic systems [70, 71, 72, 35]. The double-well character of the potential (1) requires some generalizations of this method which we present here.
In the first step we have approximated the solution of the Fokker-Planck equation (II.1) for sufficiently low biases and temperatures by a weighted sum of quasi-equilibrated sharp ($\Theta\ll 1$) Gaussian distributions [72] around the two types of local minima
@@ II.2 Full-counting statistics for the positional jump dynamics
Here $\varphi^{A}_{m}$ and $\varphi^{B}_{m}$ are the positions of the $m$-th potential minima and $P^{A}_{m}(\tau)$ and $P^{B}_{m}(\tau)$ are the corresponding time-dependent weights. These are assumed to satisfy the (Markovian) master equations (MEs)
$\frac{dP^{A}_{m}}{d\tau}$ $=$ $\sum_{j}\left(\Gamma^{A}_{2j}P^{A}_{m-j}-\Gamma^{A}_{j}P^{A}_{m}+\Gamma^{B}_{2j-1}P^{B}_{m-j}\right),$ (7)
$\frac{dP^{B}_{m}}{d\tau}$ $=$ $\sum_{j}\left(\Gamma^{B}_{2j}P^{B}_{m-j}-\Gamma^{B}_{j}P^{B}_{m}+\Gamma^{A}_{2j+1}P^{A}_{m-j}\right),$ (8)
where $\Gamma^{A}_{n}$ is the rate of a positional jump from the potential well $A$ and $\Gamma^{B}_{n}$ from the well $B$ over $n$ potential local maxima to another local minimum. Rates with even $n$ belong to positional jumps between minima of the same kind ($A\rightarrow A$, $B\rightarrow B$) and odd ones to positional jumps between minima of different kinds ($A\rightarrow B$, $B\rightarrow A$). The negative $n$’s correspond to the jump rates in the direction opposite the slope of the bias (up the hill). Following the standard FCS methodology [72], we can evaluate the $k$-dependent cumulant generating function (where $k$ is the counting field) for long times from the ME and equate it with the cumulant generating function of the full model
@@
$\mathbf{H}^{\textrm{PJ}}(k)=\left(\begin{array}[]{cc}-\sum\limits_{j}\Gamma^{A}_{j}+\Gamma^{A}_{2j}e^{ik2\pi j}&\sum\limits_{j}\Gamma^{B}_{2j-1}e^{ik(2\pi j-d)}\\ \sum\limits_{j}\Gamma^{A}_{2j+1}e^{ik(2\pi j+d)}&-\sum\limits_{j}\Gamma^{B}_{j}+\Gamma^{B}_{2j}e^{ik2\pi j}\end{array}\right).$ (20)
The $d$ factor in the exponents of the above equations is the distance between the neighboring minima $d=\varphi^{B}_{m}-\varphi^{A}_{m}$ [see Fig. 1(c)] which depends on the potential parameters $\alpha$ and $i_{b}$. In contrast to even positional jumps, where the distance traveled is always an integer multiple of $2\pi$, odd jumps overcome a distance of $2\pi j+d$ where in general $d\neq\pi$, therefore, we call these jumps fractional. Note, that we use the term fractional in its general sense of something that is less than a whole or less than a complete unit and we do not imply that the jumps are necessary rational numbers in the units of $\pi$.
Analogously to Appendix in Ref. [35] we use MCF to calculate the two counting-field-dependent eigenvalues $\lambda_{0}(k),\>\lambda_{1}(k)$ with the largest real parts of the full problem (II.1) with a modified boundary condition $W(\varphi,v;\tau\shortrightarrow\infty)=e^{i2\pi k}W(\varphi,v;\tau\shortrightarrow\infty)$ and related eigenvectors $u_{0}(k,\varphi,v)$, $u_{1}(k,\varphi,v)$  [27, Sec. 9.3]. The above two eigenvalues are in the relevant regime well separated from all the others.
@@
and $\mathbf{U}(k)$ is the square matrix of vectors $U_{0}(k)$ and $U_{1}(k)$. Note that due to the second harmonics in the potential and similarly to other non-Hermitian Hamiltonian systems [81, 82], the eigenvalues (22) can have complicated topological properties, as shown in Fig. 3. The two eigenvalues can, for small enough $\alpha$, connect at $k_{c}=1/2+j$, where $j$ is an integer, and smoothly continue each other on the Riemann surface; see Fig. 3(a). Therefore, the real parts of these two eigenvalues, plotted as functions of $k$, touch on $k_{c}$ [Fig. 3(b)], and there is a discontinuity in the imaginary part of the eigenvalues at the same points [Fig. 3(c)]. These discontinuities must be treated with care in the numerical evaluation of the eigenvalues and eigenvectors.
Figure 3: The two eigenvalues with the biggest real part plotted at the Riemann surface (a) and their projections to the real (b) and the imaginary (c) planes.
Having the reconstructed matrix $\mathbf{H}^{\textrm{PJ}}$ one can evaluate the rates of the even positional jumps, i.e., jumps between the same kind of minima with distance traveled $2\pi j=\pi n$, using the transformations
@@
and the odd ones between the different minima types, with distance traveled $\pi(n\mp 1)\pm d$, as
@@
$\langle v\rangle$ $=\lim_{\tau\rightarrow\infty}-\frac{i}{\tau}\left.\frac{\partial\mathcal{F}^{\textrm{PJ}}(k,\tau)}{\partial k}\right|_{k=0},$